# Li Tana,b , Shengrong Wanga,c

### v1

@@ 5.1 Moment bounds
×$\times\bigg{\langle}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\sigma(\overline{\mu}^{Y,N}(u)){\rm d}B_{u}^{H,i},b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\bigg{\rangle}{\rm d}s$
×$\times\bigg{\langle}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\sigma(\overline{\mu}^{Y,N}(u)){\rm d}B_{u}^{H,i},b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\bigg{\rangle}{\rm d}s$
$\leq$ $2HT^{2H-2}\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-3}{2}}|\sigma(\overline{\mu}^{Y,N}(u))|^{2}{\rm d}u$
×$\times|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))|{\rm d}s$
$\leq$ $C\Delta^{1-\alpha}\mathbb{E}\int_{0}^{t}(|Y_{\Delta}^{i,N}(s)|^{\overline{p}}+|Y_{\Delta}^{i,N}(s-\tau)|^{\overline{p}}+|\overline{Y}_{\Delta}^{i,N}(s))|^{\overline{p}}){\rm d}s$
$+C\Delta^{1-\alpha}\Delta^{(1-\alpha)\overline{p}}\mathbb{E}\int_{0}^{t}|\overline{Y}_{\Delta}^{i,N}(s))|^{\overline{p}}{\rm d}s$
$\leq$ $C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(s)|^{\overline{p}}{\rm d}s.$
To sum up, we derive
$G_{3}(t)\leq C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(s)|^{\overline{p}}{\rm d}s.$
Combining $G_{1}-G_{3}$, we get
(5.12) $\sup_{0\leq u\leq t}\mathbb{E}|\widetilde{Y}_{\Delta}^{i,N}(u)|^{\overline{p}}\leq\sup_{0\leq u\leq t}\mathbb{E}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\overline{p}/2}\leq C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}{\rm d}s.$
Noting that $|x-y|^{\overline{p}}\geq 2^{1-\overline{p}}|x|^{\overline{p}}-|y|^{\overline{p}}$, by Assumption 2.3, Lemma 5.1, we get
@@
$\geq$ $(2^{2-2\overline{p}}-4^{\overline{p}-1}\theta^{\overline{p}}K_{7}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}})|Y_{\Delta}^{i,N}(t)|^{\overline{p}}-(2^{1-\overline{p}}\lambda^{\overline{p}}+4^{\overline{p}-1}\theta^{p}K_{7}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}})|Y_{\Delta}^{i,N}(t-\tau)|^{\overline{p}}$
$-4^{\overline{p}-1}\theta^{\overline{p}}K_{7}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}}(\mathbb{W}_{q}(\mu^{Y,N}(t)),\delta_{0})^{\overline{p}}-4^{\overline{p}-1}\theta^{\overline{p}}K_{7}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}}.$
@@
$\sup_{0\leq u\leq t}\mathbb{E}|Y_{\Delta}^{i,N}(t)|^{\overline{p}}\leq(2^{2-2\overline{p}}-3\times 4^{\overline{p}-1}\theta^{\overline{p}}K_{7}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}}-2^{1-\overline{p}}\lambda^{\overline{p}})^{-1}$
×$\times\bigg{[}\sup_{0\leq u\leq t}\mathbb{E}|\widetilde{Y}_{\Delta}^{i,N}(u)|^{\overline{p}}+4^{\overline{p}-1}\theta^{\overline{p}}K_{7}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}}+(2^{1-\overline{p}}\lambda^{\overline{p}}+4^{\overline{p}-1}\theta^{\overline{p}}K_{7}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}})\mathbb{E}\parallel\xi\parallel^{\overline{p}}\bigg{]}.$
Then, we derive from (5.12),
$\sup_{0\leq u\leq t}\mathbb{E}|Y_{\Delta}^{i,N}(t)|^{\overline{p}}\leq C+C\int_{0}^{t}\sup_{0\leq u\leq t}\mathbb{E}|Y_{\Delta}^{i,N}(t)|^{\overline{p}}.$
Applying the Gronwall inequality, we complete the proof. ∎
@@ 5.2 Strong convergence rate
Lemma 5.5.
Let Assumptions 2.2-2.5 hold. For $p>1/H$, then
@@
=:absent:=:= : $J_{1}+J_{2}.$
By the Hölder inequality and Lemmas 5.1, 5.4, we get
$J_{1}\leq$ $2^{p-1}\Delta^{p-1}\mathbb{E}\int_{t_{k}}^{t_{k+1}}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))|^{p}{\rm d}s$
$\leq$ $2^{p-1}\Delta^{p-1}\mathbb{E}\int_{t_{k}}^{t_{k+1}}K_{6}\Delta^{-\alpha p}(1+|\overline{Y}_{\Delta}^{i,N}(s)|+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|+\mathbb{W}_{q}(\overline{\mu}^{Y,N}(s),\delta_{0}))^{p}{\rm d}s$
@@
Using the same technique as in (3.3), along with the Hölder inequality, the Fubini theorem and Lemmas 2.5, 5.4, we get
$J_{2}\leq C\Delta^{pH-1}\int_{t_{k}}^{t_{k+1}}\mathbb{E}(1+\mathbb{W}_{\theta}(\overline{\mu}_{s}^{Y,N},\delta_{0}))^{p}{\rm d}s\leq C\Delta^{pH}.$
Recalling that $\alpha\in(0,1/2]$, we obtian
@@
Denoting by $\widetilde{D}(t,t_{k}):=D(Y_{\Delta}^{i,N}(t-\tau))-D(Y_{\Delta}^{i,N}(t_{k}-\tau))$, and $\widetilde{b}(t,t_{k}):=b_{\Delta}(Y_{\Delta}^{i,N}(t),Y_{\Delta}^{i,N}(t-\tau),\mu^{Y,N}(t))-b_{\Delta}(Y_{\Delta}^{i,N}(t_{k}),Y_{\Delta}^{i,N}(t_{k}-\tau),\mu^{Y,N}(t_{k}))$, together with Assumption 2.3, we get
@@
For $0\leq t\leq t_{1}=\Delta$, then $t-\tau<t_{1}-\tau<0$, we derive from (5.13) and Lemma 5.4,
@@
By induction,
$\mathbb{E}\bigg{(}\sup_{t_{k}\leq t\leq t_{k+1}}|Y_{\Delta}^{i,N}(t)-Y_{\Delta}^{i,N}(t_{1})|^{p}\bigg{)}\leq$ $C\mathbb{E}\bigg{(}\sup_{t_{k}\leq t\leq t_{k+1}}|\widetilde{Y}_{\Delta}^{i,N}(t)-\widetilde{Y}_{\Delta}^{i,N}(t_{k})|^{p}\bigg{)}+C\Delta^{(1-\alpha)p}$
$\leq$ $C\Delta^{[(1-\alpha)\wedge H]p}.$
we get the desired assertion. ∎
@@
Let Assumptions 2.2-2.5 and Lemmas 5.1, 5.2 hold, then it holds that for $p>1/H$ and $\overline{p}>2(l+1)p$,
@@
For $t\in[0,T]$, denote $I(t)=\widetilde{Y}_{\Delta}^{i,N}(t)-X_{t}^{i,N}+D(X_{t-\tau}^{i,N})$, then
$I(t)=$ $I(0)+\int_{0}^{t}(b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})){\rm d}s$
@@
where $I_{0}=-\theta b_{\Delta}(\xi(0),\xi(-\tau),\delta_{0})\Delta$. The Itô formula yields,
$|I(t)|^{p}\leq$ $|I(0)|^{p}$
$+p\int_{0}^{t}|I(s)|^{p-2}\langle I(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})\rangle{\rm d}s$
$+HT^{2H-1}p(p-1)\int_{0}^{t}|I(s)|^{p-2}|\sigma(\overline{\mu}^{Y,N}(s))-\sigma(\mu_{s}^{X,N})|^{2}{\rm d}s$
$+p\int_{0}^{t}|I(s)|^{p-2}\langle I(s),\sigma(\overline{\mu}^{Y,N}(s))-\sigma(\mu_{s}^{X,N})\rangle{\rm d}B_{s}^{H,i}$
$\leq$ $|I(0)|^{p}+\sum_{i=1}^{6}H_{i}(t),$
@@
$H_{1}(t)=$ $p\int_{0}^{t}|I(s)|^{p-2}\langle I(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))$
@@
$H_{2}(t)=$ $p\int_{0}^{t}|I(s)|^{p-2}\langle I(s),b(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))$
@@
$H_{3}(t)=$ $p\int_{0}^{t}|I(s)|^{p-2}\langle I(s),b(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))$
$-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})\rangle{\rm d}s,$
$H_{4}(t)=$ $HT^{2H-1}p(p-1)\int_{0}^{t}|I(s)|^{p-2}|\sigma(\overline{\mu}^{Y,N}(s))-\sigma(\mu^{Y,N}(s))|^{2}{\rm d}s$
$H_{5}(t)=$ $HT^{2H-1}p(p-1)\int_{0}^{t}|I(s)|^{p-2}|\sigma(\mu^{Y,N}(s))-\sigma(\mu_{s}^{X,N})|^{2}{\rm d}s$
$H_{6}(t)=$ $p\int_{0}^{t}|I(s)|^{p-2}\langle I(s),\sigma(\overline{\mu}^{Y,N}(s))-\sigma(\mu_{s}^{X,N})\rangle{\rm d}B_{s}^{H,i}.$
@@
$\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}H_{1}(u)\bigg{)}$
$\leq$ $C\mathbb{E}\int_{0}^{t}|I(s)|^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))$
@@
$+C\Delta^{\alpha p}\mathbb{E}\int_{0}^{t}(1+|\overline{Y}_{\Delta}^{i,N}(s)|^{2(l+1)p}+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|^{2(l+1)p}+(\mathbb{W}_{q}(\overline{\mu}^{Y,N},\delta_{0}))^{2p}){\rm d}s$
$\leq$ $C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(s)-X_{s}^{i,N}|^{p}+C\Delta^{(1-\alpha)p}+C\Delta^{\alpha p},$
where for the last term we used the fact that $\overline{p}>2(l+1)p$. By the elementary inequality, the Young’s inequality, the Hölder inequality, Assumptions 2.3, 2.4 and Lemmas 5.1, 5.4, 5.5, we get
$\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}H_{2}(u)\bigg{)}$
$\leq$ $C\mathbb{E}\int_{0}^{t}|I(s)|^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}|b(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))-b(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))|^{p}{\rm d}s$
@@
×$\times(|\overline{Y}_{\Delta}^{i,N}(s)-Y_{\Delta}^{i,N}(s)|+|\overline{Y}_{\Delta}^{i,N}(s-\tau)-Y_{\Delta}^{i,N}(s-\tau)|+(\mathbb{W}_{q}(\overline{\mu}^{Y,N}(s),\mu^{Y,N})(s)))^{p}\bigg{]}{\rm d}s$
@@
×$\times[\mathbb{E}(|\overline{Y}_{\Delta}^{i,N}(s)-Y_{\Delta}^{i,N}(s)|+|\overline{Y}_{\Delta}^{i,N}(s-\tau)-Y_{\Delta}^{i,N}(s-\tau)|)^{2p}]^{1/2}{\rm d}s$

### v2

@@ 5.1 Moment Boundness
$\times\bigg\langle\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\sigma(\overline{\mu}^{Y,N}(u)){\rm d}B_{u}^{H,i},b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))\bigg\rangle{\rm d}s$
$\leq$ $\int_{0}^{t}\bigg\{\bigg[\mathbb{E}\bigg(\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),\sigma(\overline{\mu}^{Y,N}(u))\rangle{\rm d}B_{u}^{H,i}\bigg)^{2}\bigg]^{1/2}$
$\times\bigg[\mathbb{E}\bigg(\bigg\langle\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u)){\rm d}u,b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\bigg\rangle\bigg)^{2}\bigg]^{1/2}\bigg\}{\rm d}s$
$+$ $\int_{0}^{t}\bigg\{\bigg[\mathbb{E}\bigg(\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),\sigma(\overline{\mu}^{Y,N}(u))\rangle{\rm d}B_{u}^{H,i}\bigg)^{2}\bigg]^{1/2}$
$\times\bigg[\mathbb{E}\bigg(\bigg\langle\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\sigma(\overline{\mu}^{Y,N}(u)){\rm d}B_{u}^{H,i},b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))\bigg\rangle\bigg)^{2}\bigg]^{1/2}\bigg\}{\rm d}s$
$\leq$ $2H{\Delta}^{2H-2}\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-3}{2}}|\sigma(\overline{\mu}^{Y,N}(u))|^{2}{\rm d}u$
$\times\bigg\{\bigg(\mathbb{E}\Delta^{4-4\alpha}(1+|\overline{Y}_{\Delta}^{i,N}(s)|+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|)^{4}\bigg)^{1/4}\bigg(\mathbb{E}\Delta^{-4\alpha}(1+|\overline{Y}_{\Delta}^{i,N}(s)|+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|)^{4}\bigg)^{1/4}$
$+\bigg(\mathbb{E}\Delta^{4}(1+|\overline{Y}_{\Delta}^{i,N}(s)|)^{4}\bigg)^{1/4}\bigg(\mathbb{E}\Delta^{-4\alpha}(1+|\overline{Y}_{\Delta}^{i,N}(s)|+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|)^{4}\bigg)^{1/4}{\rm d}s$
$\leq$ $2H{\Delta}^{2H-2}\mathbb{E}\int_{0}^{t}\bigg(\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-3}{2}}|\sigma(\overline{\mu}^{Y,N}(u))|^{2}{\rm d}u\times C\Delta^{1-\alpha}\mathbb{E}\int_{0}^{t}(1+|Y_{\Delta}^{i,N}(s)|^{\overline{p}})\bigg){\rm d}s$
$\leq$ $C\Delta^{2H-1-\alpha}\int_{0}^{t}[C\Delta(1+\sup_{0\leq v\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(v)|^{\overline{p}})]\times[C(1+\sup_{0\leq v\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(v)|^{\overline{p}})]\mathrm{d}s$
$=$ $C\Delta^{2H-\alpha}\int_{0}^{t}(1+\sup_{0\leq v\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(v)|^{\overline{p}})^{2}\mathrm{d}s$
$\leq$ $C\Delta^{2H-1-\alpha}(1+\sup_{0\leq v\leq t}\mathbb{E}|Y_{\Delta}^{i,N}(v)|^{\overline{p}})\int_{0}^{t}(1+\sup_{0\leq v\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(v)|^{\overline{p}})\mathrm{d}s$
To sum up, we conclude
$G_{3}(t)\leq C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}{\rm d}s+C\int_{0}^{t}(t-s)^{2H-1}\bigg(\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}\bigg){\rm d}s.$
Combining $G_{1}(t)-G_{3}(t)$ together, it shows
(5.12) $\sup_{0\leq u\leq t}\mathbb{E}|\widetilde{Y}_{\Delta}^{i,N}(u)|^{\overline{p}}\leq$ $\sup_{0\leq u\leq t}\mathbb{E}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\overline{p}/2}\leq C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}{\rm d}s$
$+C\int_{0}^{t}(t-s)^{2H-1}\bigg(\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}\bigg){\rm d}s.$
Noting that $|x-y|^{\overline{p}}\geq 2^{1-\overline{p}}|x|^{\overline{p}}-|y|^{\overline{p}}$, Assumption 2.3 and Lemma 5.1 give
@@
$\geq$ $(2^{2-2\overline{p}}-4^{\overline{p}-1}\theta^{\overline{p}}K_{6}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}})|Y_{\Delta}^{i,N}(t)|^{\overline{p}}-(2^{1-\overline{p}}\lambda^{\overline{p}}+4^{\overline{p}-1}\theta^{p}K_{6}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}})|Y_{\Delta}^{i,N}(t-\tau)|^{\overline{p}}$
$-4^{\overline{p}-1}\theta^{\overline{p}}K_{6}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}}\mathbb{W}_{2}^{\overline{p}}(\mu^{Y,N}(t),\delta_{0})-4^{\overline{p}-1}\theta^{\overline{p}}K_{6}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}}.$
@@
$\sup_{0\leq u\leq t}\mathbb{E}|Y_{\Delta}^{i,N}(t)|^{\overline{p}}\leq(2^{2-2\overline{p}}-3\times 4^{\overline{p}-1}\theta^{\overline{p}}K_{6}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}}-2^{1-\overline{p}}\lambda^{\overline{p}})^{-1}$
$\times\bigg[\sup_{0\leq u\leq t}\mathbb{E}|\widetilde{Y}_{\Delta}^{i,N}(u)|^{\overline{p}}+4^{\overline{p}-1}\theta^{\overline{p}}K_{6}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}}+(2^{1-\overline{p}}\lambda^{\overline{p}}+4^{\overline{p}-1}\theta^{\overline{p}}K_{6}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}})\mathbb{E}\|\xi\|^{\overline{p}}\bigg].$
Further, it follows from (5.12) that
$\sup_{0\leq u\leq t}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}\leq C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}{\rm d}s+C\int_{0}^{t}(t-s)^{2H-1}\bigg(\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}\bigg){\rm d}s.$
Consequently, the desired result can be shown by applying Lemma 2.8. This completes the proof. ∎
@@ 5.2 Strong Convergence Rate
Lemma $5.$
Let Assumptions 2.2-2.5 hold. For $p\geq 2$, then
@@
$:=$ $W_{1}(t)+W_{2}(t).$
With the Hölder inequality and Lemmas 5.1, 5.4, it follows
$W_{1}(t)=$ $2^{p-1}\mathbb{E}\bigg(\sup_{t_{k}\leq t\leq t_{k+1}}\bigg\lvert\int_{t_{k}}^{t}b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s)){\rm d}s\bigg\rvert^{p}\bigg)$
$\leq$ $2^{p-1}\Delta^{p-1}\mathbb{E}\int_{t_{k}}^{t_{k+1}}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))|^{p}{\rm d}s$
$\leq$ $2^{p-1}\Delta^{p-1}\mathbb{E}\int_{t_{k}}^{t_{k+1}}K_{6}\Delta^{-\alpha p}(1+|\overline{Y}_{\Delta}^{i,N}(s)|+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|+\mathbb{W}_{2}^{p}(\overline{\mu}^{Y,N}(s),\delta_{0})){\rm d}s$
@@
Using the same techniques as the proof process of [15, Theorem 3.1], along with the Hölder inequality, the Fubini theorem and Lemmas 2.5, 5.4, we derive
$W_{2}(t)=$ $2^{p-1}\mathbb{E}\bigg(\sup_{t_{k}\leq t\leq t_{k+1}}\bigg\lvert\int_{t_{k}}^{t}\sigma(\overline{\mu}^{Y,N}(s)){\rm d}B_{s}^{H,i}\bigg\rvert^{p}\bigg)$
$\leq$ $C\Delta^{pH-1}\mathbb{E}\int_{t_{k}}^{t_{k+1}}(1+\mathbb{W}_{2}(\overline{\mu}_{s}^{Y,N},\delta_{0}))^{p}{\rm d}s\leq C\Delta^{pH}.$
Recalling that $\alpha\in(0,1/2]$, then we conclude
@@
Denote $\widetilde{D}(t,t_{k}):=D(Y_{\Delta}^{i,N}(t-\tau))-D(Y_{\Delta}^{i,N}(t_{k}-\tau))$, and $\widetilde{b}(t,t_{k}):=b_{\Delta}(Y_{\Delta}^{i,N}(t),Y_{\Delta}^{i,N}(t-\tau),\mu^{Y,N}(t))-b_{\Delta}(Y_{\Delta}^{i,N}(t_{k}),Y_{\Delta}^{i,N}(t_{k}-\tau),\mu^{Y,N}(t_{k}))$, together with Assumption 2.3
@@
For $0\leq t<t_{1}=\Delta$, then $t-\tau<t_{1}-\tau<0$, we derive from (5.13) and Lemma 5.4,
@@
By induction, the desired assertion can be shown. ∎
@@
Let Assumptions 2.2-2.5 hold. Fix any $\overline{p}>4(l+1)$ in Lemma 5.4. Then it holds that for $2\leq p<\overline{p}/2(l+1)$,
@@
For $t\in[0,T]$, denote $\Lambda(t)=\widetilde{Y}_{\Delta}^{i,N}(t)-X_{t}^{i,N}+D(X_{t-\tau}^{i,N})$, then
$\Lambda(t)=$ $\Lambda(0)+\int_{0}^{t}(b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})){\rm d}s$
@@
where $\Lambda({0})=-\theta b_{\Delta}(\xi(0),\xi(-\tau),\delta_{0})\Delta$. The fractional Itô formula yields,
$|\Lambda(t)|^{p}\leq$ $|\Lambda(0)|^{p}$
$+p\int_{0}^{t}|\Lambda(s)|^{p-2}\langle\Lambda(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})\rangle{\rm d}s$
$+H(2H-1)p(p-1)\int_{0}^{t}|\Lambda(s)|^{p-2}\int_{0}^{s}(s-r)^{2H-2}\|\sigma(\overline{\mu}^{Y,N}(r))-\sigma(\mu_{r}^{X,N})\|^{2}{\rm d}r{\rm d}s$
$+p\int_{0}^{t}|\Lambda(s)|^{p-2}\langle\Lambda(s),\sigma(\overline{\mu}^{Y,N}(s))-\sigma(\mu_{s}^{X,N})\rangle{\rm d}B_{s}^{H,i}$
$\leq$ $|\Lambda(0)|^{p}+\sum_{i=1}^{6}Z_{i}(t),$
@@
$Z_{1}(t)=$ $p\int_{0}^{t}|\Lambda(s)|^{p-2}\langle\Lambda(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))$
@@
$Z_{2}(t)=$ $p\int_{0}^{t}|\Lambda(s)|^{p-2}\langle\Lambda(s),b(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))$
@@
$Z_{3}(t)=$ $p\int_{0}^{t}|\Lambda(s)|^{p-2}\langle\Lambda(s),b(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})\rangle{\rm d}s,$
$Z_{4}(t)=$ $H(2H-1)p(p-1)\int_{0}^{t}|\Lambda(s)|^{p-2}\int_{0}^{s}(s-r)^{2H-2}\|\sigma(\overline{\mu}^{Y,N}(r))-\sigma(\mu^{Y,N}(r))\|^{2}{\rm d}r{\rm d}s$
$Z_{5}(t)=$ $H(2H-1)p(p-1)\int_{0}^{t}|\Lambda(s)|^{p-2}\int_{0}^{s}(s-r)^{2H-2}\|\sigma(\mu^{Y,N}(r))-\sigma(\mu_{r}^{X,N})\|^{2}{\rm d}r{\rm d}s$
$Z_{6}(t)=$ $p\int_{0}^{t}|\Lambda(s)|^{p-2}\langle\Lambda(s),\sigma(\overline{\mu}^{Y,N}(s))-\sigma(\mu_{s}^{X,N})\rangle{\rm d}B_{s}^{H,i}.$
@@
$\mathbb{E}\bigg(\sup_{0\leq u\leq t}Z_{1}(u)\bigg)$
$\leq$ $C\mathbb{E}\int_{0}^{t}|\Lambda(s)|^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))$
@@
$+C\Delta^{\alpha p}\mathbb{E}\int_{0}^{t}(1+|\overline{Y}_{\Delta}^{i,N}(s)|^{2(l+1)p}+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|^{2(l+1)p}+\mathbb{W}_{2}^{2p}(\overline{\mu}^{Y,N},\delta_{0})){\rm d}s$
$\leq$ $C\int_{0}^{t}\mathbb{E}\left(\sup_{0\leq u\leq s}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}\right){\rm d}s+C\Delta^{(1-\alpha)p}+C\Delta^{\alpha p},$
where for the last term we used the fact that $2(l+1)p<\overline{p}$. Combining the elementary inequality, the Young’s inequality, the Hölder inequality, Assumptions 2.3, 2.4 and Lemmas 5.1, 5.4, 5.5, it follows
$\mathbb{E}\bigg(\sup_{0\leq u\leq t}Z_{2}(u)\bigg)$
$\leq$ $C\mathbb{E}\int_{0}^{t}|\Lambda(s)|^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}|b(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))-b(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))|^{p}{\rm d}s$
@@
$\times(|\overline{Y}_{\Delta}^{i,N}(s)-Y_{\Delta}^{i,N}(s)|+|\overline{Y}_{\Delta}^{i,N}(s-\tau)-Y_{\Delta}^{i,N}(s-\tau)|+\mathbb{W}_{2}^{p}(\overline{\mu}^{Y,N}(s),\mu^{Y,N})(s))\bigg]{\rm d}s$
@@
$\times[\mathbb{E}(|\overline{Y}_{\Delta}^{i,N}(s)-Y_{\Delta}^{i,N}(s)|+|\overline{Y}_{\Delta}^{i,N}(s-\tau)-Y_{\Delta}^{i,N}(s-\tau)|)^{2p}]^{1/2}{\rm d}s$